# Diffusion-Based mmWave Radar Point Cloud Enhancement Driven by Range Images

### v1

@@ IV-C2 Training Objective
where $\lambda_{m},\lambda_{p},\lambda_{c}$ denotes the weight coefficients of each loss function, used to balance multi-objective optimization during model training
@@ V Experiments and Results
In this section, we compare the proposed method against baseline methods on both the public dataset (Coloradar) and the self-constructed dataset to validate the superior performance of our approach.
@@ V-A Benchmark Comparisons
We performed benchmark comparisons with five baseline methods, including RPDNet [5], SO-CFAR [9], CA-CFAR [10], GO-CFAR [11], and OS-CFAR [12]. All of these methods, like the one we propose, are capable of generating authentic 3D point clouds. We excluded comparisons with the diffusion model-based methods proposed by Zhang et al. [7] and Luan et al. [8], as these methods are solely designed to generate high-quality BEV maps.
V-A1 Dataset
The ColoRadar dataset provides LiDAR point cloud data and raw mmWave ADC data for seven different environments, including indoor scenes (Arpg Lab, Ec Hallways, and Aspen Corporation), outdoor scenes (Longboard and Outdoor), and mining environments (Edgar Army and Edgar Classroom). Each environment contains multiple trajectories. We use the first three trajectories of each scene as the training set, while the remaining trajectories are used as the test set. In addition, using the handheld data collection platform shown in Fig. 7, we respectively collected four trajectories in the hall and corridor of the Robotics Institute of Zhejiang University, totaling 6,636 frames. Similarly, we set the first three trajectories of each scene as the training set, and the others as the testing set.
@@ V-A1 Dataset
V-A2 Data Collection Platform
We customized a handheld data collection platform, as shown in Fig. 7. It consists of an NVIDIA Jetson Orin NX for recording sensor data, a forward-tilted Livox Mid-360 LiDAR, and a BMI270 IMU to estimate real-time platform poses with the assistance of Fast-LIO [21] for visualization. Additionally, it is equipped with a TI-AWR1843BOOST and TI-DCA1000EVM for collecting mmWave radar raw data, and an Intel RealSense Depth Camera D435 to capture image data for visualization.
V-A3 Preprocessing
Additionally, considering that mmWave radar is not sensitive to low-reflectivity objects, we first employed DBSCAN to cluster the combined point cloud of mmWave radar and LiDAR. Then, We applied additional filtering to the LiDAR point cloud using labels from mmWave radar to eliminate objects that are difficult for the mmWave radar to detect. This step ensures that the neural network does not learn features undetectable by mmWave radar, thereby preventing the neural network from being misled during training.
V-A4 Implementation
In the process of mmWave point cloud processing, we employed the mmWave radar signal processing techniques described in Section III, alongside the CFAR detector, to acquire the raw mmWave radar point cloud data. The performance of the CFAR detector is notably influenced by the threshold setting. A higher threshold enables superior denoising performance; however, it also leads to a loss of information. Under such conditions, it becomes nearly impossible to match the density of the mmWave radar point cloud to that of LiDAR point clouds. Therefore, to preserve as much raw mmWave radar information as possible, we configured the CFAR detector’s threshold to approach zero.
@@ V-A4 Implementation
V-B Qualitative Comparison
The mmWave radar point clouds generated by both the proposed and baseline methods, alongside the ground truth LiDAR point clouds, are presented in Fig. 1 and Fig. 5. The results demonstrate that our method outperforms all baseline methods in terms of both denoising capability and densification across all scenes. Our method achieves remarkable reconstruction of the curved features of rock walls in mining environments, as well as the sharp corner features in intricate environments. Among the baseline methods, OS-CFAR [12], a traditional technique for processing mmWave data, estimates noise levels by sorting the signal intensity within the observation window to adjust the detection threshold. However, it is highly susceptible to outliers, leading to poor detection performance. In contrast, RPDNet [5] utilizes LiDAR point clouds to assist in training a GAN, thereby enabling it to eliminate more clutter and detect more targets. Nevertheless, RPDNet [5] still relies on traditional DOA estimation methods, resulting in outputs that contain considerable noise.
V-C Quantitative Comparison
@@
We begin by presenting the evaluation metrics employed in our quantitative comparison: Chamfer Distance (CD), Modified Hausdorff Distance (MHD), and F-Score. CD quantifies the global similarity between two point clouds by summing the distances from each point to the nearest point in the other point cloud. MHD evaluates the maximum disparity between two point clouds by calculating their forward and backward Hausdorff distances. F-Score, defined as the harmonic mean of precision $P$ and recall $R$, is utilized to assess the quality of point cloud reconstruction. The results of the quantitative analysis are provided in Tab.I and Tab.II, while Fig. 6 illustrates the cumulative distribution function (CDF) curves of the Chamfer Distance.
The results demonstrate that the proposed method significantly outperforms the baseline approaches across all scenes in both datasets. The RPDNet [5] method is only marginally comparable to the proposed approach concerning the CD metric in the Outdoor scenes of the ColoRadar dataset and the MHD metric in the Corridor scenes of the self-constructed dataset. This arises from the inherent complexity of the Outdoor scene, which is populated by numerous targets exhibiting low reflectivity and small sizes, thus hindering the ability of mmWave radar to effectively observe these objects, thereby affecting the model’s performance. Furthermore, The Corridor scene in the self-constructed dataset contains a significant amount of glass. The smooth texture of glass has the potential to deflect millimeter waves in various directions, generating virtual images, while the multi-layered glass structures may cause multiple reflections of the radar signals, resulting in interference that detracts from the efficacy of the proposed method.
@@ VI Conclusion
This paper proposes a novel approach for generating high-quality mmWave radar point clouds. By fully leveraging the powerful generation capability of the diffusion model, the proposed method can effectively process the sparse and noisy original mmWave radar point cloud of a single frame and obtain an accurate and dense point cloud similar to LiDAR. Our approach employs range images aligned with human perception as proxy representations for point clouds and combines them with image diffusion models for the first time to solve this task. Owing to the projection that aligns with human perception, the range image representation closely resembles natural images, thereby facilitating the transfer of knowledge from the pre-trained image diffusion model and substantially enhancing overall performance. The proposed method is validated on public and self-constructed datasets, significantly outperforming baseline methods in terms of both point cloud quality and density.

### v2

@@ IV-C2 Training Objective
where $\lambda_{m},\lambda_{p},\lambda_{c}$ denote the weight coefficients of each loss function, used to balance multi-objective optimization during model training
@@ V Experiments and Results
In this section, we conduct benchmark comparisons, real-world experiments, and ablation experiments on both the public dataset (Coloradar) and the self-constructed dataset to validate the superior performance of our method.
@@ V-A Benchmark Comparisons
We performed benchmark comparisons with four baseline methods, including OS-CFAR [13], MUSIC [14], RPDNet [5], and DenserRadar [15].
V-A1 Dataset Configuration
The ColoRadar dataset provides LiDAR point cloud data and raw mmWave ADC data for seven different environments. Each environment contains multiple trajectories. We employ the data from the four scenes enumerated in Table I for both training and testing. The first three trajectories of each scene constitute the training set, while the remaining trajectories are designated as the test set. To ensure a fair comparison, the dataset configurations of RPDNet [5] and DenserRadar [15] methods are aligned with those adopted in this study.
V-A2 Dataset Preprocessing
Additionally, considering that mmWave radar is not sensitive to low-reflectivity objects, we first employed DBSCAN to cluster the combined point cloud of mmWave radar and LiDAR. Then, we applied additional filtering to the LiDAR point cloud using labels from the mmWave radar to eliminate objects that are difficult for the mmWave radar to detect. This step ensures that the neural network does not learn features undetectable by mmWave radar, thereby preventing the neural network from being misled during training.
@@ V-A2 Dataset Preprocessing
V-A3 Implementation
In the process of mmWave point cloud processing, we obtain the original mmWave radar point cloud data based on the CFAR. The performance of the CFAR detector is notably influenced by the threshold setting. A higher threshold enables superior denoising performance; however, it also leads to a loss of information. Under such conditions, it becomes nearly impossible to match the density of the mmWave radar point cloud to that of LiDAR point clouds. Therefore, to preserve as much raw mmWave radar information as possible, we configured the CFAR detector’s threshold to approach zero.
@@ V-A3 Implementation
V-A4 Qualitative Comparison
The mmWave radar point clouds generated by both the proposed and baseline methods, alongside the ground truth LiDAR point clouds, are presented in Fig. 1 and Fig. 4. The results demonstrate that our method outperforms all baseline methods in terms of density across all scenes. Our approach produces high-fidelity scene details, reconstructs the curved geometry of objects, and maintains sharp corner features under complex environmental conditions. In contrast to the OS-CFAR [13] and MUSIC [14], both RPDNet [5] and DenserRadar [15] employ deep networks as detectors and achieve significantly improved detection results. Nevertheless, the point clouds produced by these methods remain less dense and accurate than those generated by our proposed approach. Furthermore, DenserRadar [15] operates directly on the voluminous raw radar data in the form of a 4-dimensional tensor (cube), which necessitates a model with substantially greater capacity.
V-A5 Quantitative Comparison
@@
We begin by presenting the evaluation metrics employed in our quantitative comparison: Chamfer Distance (CD), Modified Hausdorff Distance (MHD), and F-Score. CD quantifies the global similarity between two point clouds by summing the distances from each point to the nearest point in the other point cloud. MHD evaluates the disparity between two point clouds by calculating their forward and backward Hausdorff distances. F-score, defined as the harmonic mean of precision $P$ and recall $R$, is utilized to assess the quality of point cloud reconstruction. The results of the quantitative analysis are provided in Tab.I, while Fig. 7 illustrates the cumulative distribution function (CDF) curves of the Chamfer Distance.
The results demonstrate that the proposed method significantly outperforms the baseline approaches across all scenes in the ColoRadar datasets. In the Edgar Mine scene of ColoRadar datasets, the RPDNet [5] methodis only slightly better than our approach with respect to the MHD metric. This is primarily because the Edgar Mine scene is dominated by walls with simple semantics and geometry, which limits the advantage of our method that leverages rich semantic and geometric information.
TABLE III: Ablation Experiment of the Data Representation Format.
  Method Hallways Edgar Mine Aspen Lab Outdoor
CD(m) $\downarrow$ MHD(m) $\downarrow$ F-Score($\%$) $\uparrow$ CD(m) $\downarrow$ MHD(m) $\downarrow$ F-Score($\%$) $\uparrow$ CD(m) $\downarrow$ MHD(m) $\downarrow$ F-Score($\%$) $\uparrow$ CD(m) $\downarrow$ MHD(m) $\downarrow$ F-Score($\%$) $\uparrow$
  Ours-BEV 2.452 0.548 51.808 5.127 1.154 30.374 2.462 0.723 47.234 20.155 8.358 18.113
Ours-RI 2.259 0.522 48.6 2.754 0.644 36.3 2.042 0.676 48.8 6.615 1.948 26.0
Note: Ours-RI denotes our method using Range Image as the data representation format, while Ours-BEV represents our approach with Bird’s Eye View representation.
(a) Hallways
(b) Edgar Mine
(c) Aspen
(d) Outdoors
Figure 6: Ablation study on the number of channels in mmWave radar range images. The y-axis denotes the relative proportion of each metric with respect to the case when the number of channels is one.
V-B Ablation Experiments
Beyond benchmark comparisons, we conducted ablation experiments to substantiate the necessity of the proposed method.
V-B1 Data Representation Format
In this experiment, the data representation in our method is switched from range images to BEV, a format widely used in prior studies [7, 8, 9], while all other components remain unchanged. Table III reports the quantitative results: representing the data as range images yields substantially higher point‑cloud quality than BEV. This suggests that range images, which align with human perceptual priors, more effectively exploit the capabilities of the diffusion model, thereby underscoring the efficacy of the proposed approach.
V-B2 Range Image Channel Number
Furthermore, we performed an ablation study on the range image channels. Only the number of range image channels was altered in this experiment, and all other settings remained identical to the benchmark comparisons. The results are presented in Fig. 6. It is immediately noticeable that all multi-channel range images substantially surpassed their single-channel counterparts in performance. This phenomenon stems from the intrinsic penetrative capability of mmWave radar, which produces numerous data points along identical trajectories. Employing solely single-channel range imagery culminates in multiple points being superimposed upon identical pixels, thereby precipitating substantial information loss. The experiment substantiates the imperative for employing multi-channel range images, with 16-channel configurations yielding better performance.
(a) Hallways
(b) Edgar Mine
(c) Aspen
(d) Outdoors
V-C Real-world Experiments
We further assess the generalization capability of the proposed method by employing our self-constructed datasets.
V-C1 Data Collection Platform
We customized a handheld data collection platform, as shown in Fig. 5. It consists of an NVIDIA Jetson Orin NX for recording sensor data, a Livox Mid-360 LiDAR, and a BMI270 IMU to estimate real-time platform poses with the assistance of Fast-LIO [23]. Additionally, it is equipped with a TI-AWR1843BOOST and TI-DCA1000EVM for collecting mmWave radar raw data, and an Intel RealSense D435 to capture image data for visualization.Using the handheld data collection platform, we respectively collected four trajectories in the hall and corridor, totaling 6,636 frames.
V-C2 Generalization Capability Tests
We deployed the model trained on the ColoRadar dataset directly on our self-collected dataset, without any further fine-tuning, and benchmarked it against CFAR-based [13] baselines. This test has achieved robust generalization across heterogeneous scenes, LiDAR sensors, and radar parameterizations. Quantitative results are reported in Table II, with metrics computed using LiDAR point clouds as ground truth. Even under unseen scenarios and sensor configurations, the proposed method substantially outperforms conventional CFAR-based approaches, compellingly attesting to its efficacy and generalization capacity.
@@ VI Conclusion
This paper proposes a novel approach for generating high-quality mmWave radar point clouds. By fully leveraging the powerful generation capability of the diffusion model, the proposed method can effectively process the sparse and noisy original mmWave radar point cloud of a single frame and obtain a point cloud similar to LiDAR. Our approach employs range images aligned with human perception as proxy representations for point clouds and combines them with image diffusion models for the first time to solve this task. Owing to the projection that aligns with human perception, the range image representation closely resembles natural images, thereby facilitating the transfer of knowledge from the pre-trained image diffusion model and substantially enhancing overall performance. The proposed method is validated on public and self-constructed datasets, significantly outperforming baseline methods in terms of both point cloud quality and density.